OpenStudy (anonymous):

solve this system using substitution 4y=5x-1 3x-2y=1

OpenStudy (the_fizicx99):

4y=5x-1 /4 /4 /4 y = 5/4x - 0.25 Plug in : y = 5/4x - 0.25 in the equation 3x-2y=1 for y

OpenStudy (the_fizicx99):

3x-2(1.25x -0.25)=1

OpenStudy (anonymous):

Pexplain better plz